Taylor Larson is paid on the first day of the month and $80 is automatically deducted from her pay and deposited in a savings account. If the account pays 2.5% interest compounded monthly, how much will be in the account after 3 years and 9 months? $3777.89
1
Expert's answer
2020-04-14T15:07:28-0400

1."S=P(1+i\times n)=80(1+0.025)=82" in month

2."82\times45=3690" for 45 months

3."80\times(1+\frac{0.025}{12})^{45}=87.85" for 45 months

4.3690+87.85=3777.85